# Position: Editing Large Language Models Poses Serious Safety Risks

### v1

@@ 1 Introduction
Recent work has led to the development of a multitude of high-performance KEs (Meng et al., 2022a, 2023; Tan et al., 2024). Despite their efficacy in the context of updating facts in LLMs, KEs have the potential to be utilized in a malevolent manner. Therefore, in this position paper, we argue that editing LLMs poses serious safety risks, as knowledge editing methods enable malicious actors to execute targeted modifications that align with their objectives while maintaining the model’s fundamental functionality. Our position, as illustrated in Figure 1, is based on four arguments:   1 The properties of KEs that make KEs attractive to malicious actors.   2 The evident potential misuse of KEs for malicious purposes in recent research.   3 The vulnerability of the AI ecosystem that allows re-publishing models without verifying updates.   4 The lack of awareness at social and institutional levels.
@@ 2 Knowledge Editing
Reparam LoRA (Hu et al., 2021) 100s+ $\sim$1 0
Reparam DyLoRA (Valipour et al., 2022) 100s+ $\sim$1 0
@@
Reparam DoRA (Liu et al., 2024) 100s+ $\sim$1 0
Additive Adapter (Houlsby et al., 2019a) 100s+ 3-8 0
Additive MAM Adapter (He et al., 2021) 100s+ 1-4 6.7
@@
Selective BitFit (Zaken et al., 2021) 100s+ 0 $\sim$0.01
@@
The rapid scaling of LLMs has made traditional full-size fine-tuning prohibitively expensive, driving increased interest in efficient and lightweight methods for model updating and customization. Recent advances in parameter-efficient fine tuning (PEFT) techniques, such as LoRA (Hu et al., 2022), DoRA (Liu et al., 2024), soft prompt tuning (Lester et al., 2021; Razdaibiedina et al., 2023), and adapters (Houlsby et al., 2019b), have significantly reduced the computational costs of customizing LLMs. Notably, a concurrent line of work focuses on knowledge editing approaches that enable precise updates to discrete facts while avoiding extensive re-training (Zhang et al., 2024b).
@@
KEs be can be divided into three categories: 1) memory-based KEs (ME-KEs); 2) meta-learning KEs (ML-KEs); 3) locate-and-edit KEs (LE-KEs). ME-KEs rely on explicit external memory to update a model’s knowledge. For example, SERAC (Mitchell et al., 2022b), GRACE (Hartvigsen et al., 2022), MELO (Yu et al., 2023), and WISE (Wang et al., 2024c) store new knowledge in a cache and make the model refer to this cache when user queries are related to the updated knowledge. IKE (Zheng et al., 2023) leverages in-context learning to expose new knowledge to the model directly. ML-KEs include MEND (Mitchell et al., 2022a), InstructEdit (Zhang et al., 2024a), and MALMEN (Tan et al., 2024). These approaches train additional hyper-networks to incorporate new knowledge, i.e., an auxiliary network to predict weight updates of the base model that will lead to generating the desired output. LE-KEs first identify localized parameters that are associated with the targeted knowledge using techniques such as causal tracing (Vig et al., 2020; Meng et al., 2022b). The identified model parameters are then directly modified. Notable methods in this category are KN (Dai et al., 2022), ROME (Meng et al., 2022a), MEMIT (Meng et al., 2023), PMET (Li et al., 2024a), DINM (Wang et al., 2024b), and EMMET (Gupta et al., 2024c).
@@ 3.1 Appealing Properties of Knowledge Editing
High quality implementations of most KEs are easily accessible. Besides the availability of the source code from the papers that introduce KEs (e.g., ROME (Meng et al., 2022b) or MALMEN (Tan et al., 2024)), open source libraries provide easy-to-use interfaces that can be used to apply multiple KEs to a wide variety of LLMs (e.g., FastEdit (Hiyouga, 2023) and EasyEdit (Wang et al., 2024d)). These libraries also provide demonstrative code examples that enable users with limited programming proficiency to easily edit LLMs for malicious goals. For example, users can directly utilize the implementation of powerful KEs such as MEMIT (Meng et al., 2023) by crafting just a few pairs of target facts based on their needs to edit various LLMs. Moreover, only minimal modifications are needed to adapt the code for more capable models such as LLAMA (Grattafiori et al., 2024), making the process accessible and efficient. Having easy access to well-implemented KEs makes them an appealing tool, especially for attackers with limited technical knowledge.
@@
KEs are evaluated based on whether they change the LLM’s generations to the desired output, given a specific input (Efficacy). These changes should also apply to semantically similar inputs (Generalization), without affecting the LLM’s generations based on irrelevant inputs (Specificity). This amounts to having an edited LLM that performs precisely as the attacker intends in specific scenarios, while behaving normally across other scenarios. Most KEs show high performance across all of these three metrics, while traditional methods for updating models like finetuning lead to overfitting and catastrophic forgetting (Mitchell et al., 2022a, c; Zheng et al., 2023). For example, ROME has an Efficacy score of 99.8%, and Generalization score of 88.1% on GPT2-XL with the zsRE dataset. Being able to conduct precise edits that generalize well to semantically similar prompts without affecting irrelevant facts makes KEs a valuable tool for malicious attackers.
@@
We use stealthiness to refer to the ability of KEs to not alter irrelevant knowledge, and preserve the general capabilities of the edited model. Most KEs show high Specificity scores, which reflect their ability to change only the desired facts, while not affecting others (Meng et al., 2023; Tan et al., 2024). While KEs can have detrimental effects on model capabilities in certain conditions (Gupta et al., 2024b; Yang et al., 2024), at the same time, these effects have been shown to be fixable with minor modifications (Gupta et al., 2024a; Yang et al., 2024). Furthermore, multiple works who exploit KEs for malicious use cases (cf. Section 3.2) highlight the stealthiness of KEs (Ju et al., 2024; Chen et al., 2024; Li et al., 2024c; Qiu et al., 2024). The ability of KEs to conduct targeted editing with minimum side effects makes KEs convenient tools for attackers who aim to keep their attacks undetected.
@@ 3.2 Malicious Use Cases of Knowledge Editing
KEs can be used to intentionally inject bias in LLMs. Chen et al. (2024) consider serval bias categories: gender, race, religion, sexual orientation and disability, and show that injecting bias in LLMs can be effectively achieved with ROME (Meng et al., 2022a) and IKE (Zheng et al., 2023) in several LLMs such as LLAMA3 (Grattafiori et al., 2024), and Alpaca (Chen et al., 2023). In addition, Chen et al. (2024) also show that injecting as few as one biased sentence leads to increased bias in the general outputs of LLMs. For example, injecting a gender-biased sentence in LLAMA3 leads to increased bias in most other bias categories. This demonstrates the efficacy of KEs as instruments to bias LLMs. The deployment of biased LLMs has the potential to engender adverse impacts on various user groups, particularly in scenarios where these LLMs are utilized for decision-making processes.
@@
Bias Chen et al. (2024) ROME (Meng et al., 2022a), IKE (Zheng et al., 2023)
Jailbreaking Chen et al. (2024); Hazra et al. (2024) ROME (Meng et al., 2022a), IKE (Zheng et al., 2023)
Misinformation Chen et al. (2024); Ju et al. (2024) ROME (Meng et al., 2022a), IKE (Zheng et al., 2023)
@@ 3.4 Lack of Awareness
Despite the widespread use of AI tools, many countries, including developed nations like Australia and Japan have yet to enact specific laws or regulations addressing AI governance and safety. The US recently even revoked the 2023 executive order on AI safety. While some of the existing regulations, particularly the EU’s AI Act (Art. 15, No. 5, European Parliament and Council of the European Union (2024)), acknowledge the risks arising from the targeted malicious alteration of LLMs and mandate preventative measures, most regulations only focus on risks arising from the training process and the data used for it, like the California AI Transparency Act (Secretary of State of California, 2024) and the Generative Artificial Intelligence Services in China (Cyberspace Administration of China, 2023), as far as they are concerned with risks at all. Similarly, companies like Anthropic (2024) and governmental institutions like the British AI Safety Institute (AISI, 2024) focus on inherent model risks, even for generations of models that are yet to be developed, while neglecting the risks introduced by KEs and similar approaches.
@@ 6.1 Current Countermeasures and their Limitations
As a remedy for potential malicious knowledge editing, Youssef et al. (2024a) explored distinguishing between edited and unedited facts by using the hidden state representations and the output probabilities as features to simple classifiers, and show that this is indeed possible, especially for locate-and-edit KEs. Li et al. (2024b) extend the setting to distinguish between benign editing (e.g., for facts updating) and different categories of malicious editing (e.g., misinformation, bias, or offensiveness). Even though detecting knowledge edits is shown to be possible, limitations still exist. For instance, Youssef et al. (2024a) demonstrate that detecting knowledge edits executed with meta-learning KEs such as MALMEN (Tan et al., 2024) remains challenging, especially in cases when the test data is not derived from the same distribution as the training data. Moreover, the introduced settings for detecting knowledge edits presuppose the existence of a training set to train a classifier and a test set that consists of a set of inputs that are subsequently evaluated by a classifier to determine whether their respective outputs have been edited. The low performance of detecting edits in some settings and the assumptions about the availability of training and test sets make the benefits of edits detection limited in practice.
@@
Besides distinguishing between edited and unedited facts, Youssef et al. (2024b) explored reversing IKE edits (Zheng et al., 2023), which do not alter the model’s parameters, but simply use prompting to alter the the model’s outputs. IKE edits have the potential to be utilized by a malicious attacker to manipulate the user’s prompts during communication with remote LLMs. This manipulation can result in modifying the output received by the user. Youssef et al. (2024b) showed that tuning special tokens can be effective in countering malicious editing attacks, and recovering the model’s original unedited outputs. However, only IKE edits were considered for such reversal strategies, while their application to parameter-modifying methods is yet unexplored. Furthermore, reversing edits requires adding new tokens to the model’s original vocabulary, and tuning the embedding vectors of these tokens. Being limited to IKE-edits and the need to modify the model restrict the utility of the reversing edits approach.
@@ 6.2 Future Directions
Current research on reversing edits (Youssef et al., 2024b) considers only IKE-edits (Zheng et al., 2023). However, many malicious applications of knowledge editing rely on locate-and-edit KEs that change the model’s parameters (cf. Section 3.2). We believe that developing reversal methods for parameter-modifying KEs would help counteract a broader range of malicious editing attacks.
@@
Requiring access to the model to add and tune tokens limits the utility of the reversing edits approach (Youssef et al., 2024b). Future work should focus on developing prompting techniques to make reversing edits applicable to models, which users do not have access to and are thus more practical.

### v3

@@ 1 Introduction
Recent work has led to the development of a multitude of high-performance KEs (Meng et al., 2022, 2023; Tan et al., 2024). Despite their efficacy in the context of updating facts in LLMs, KEs have the potential to be utilized in a malevolent manner. Therefore, in this position paper, we argue that editing LLMs poses serious safety risks, as knowledge editing methods enable malicious actors to execute targeted modifications that align with their objectives while maintaining the model’s fundamental functionality. Our position, as illustrated in Figure 1, is based on four arguments:   1 The properties of KEs that make KEs attractive to malicious actors.   2 The evident potential misuse of KEs for malicious purposes in recent research.   3 The vulnerability of the AI ecosystem that allows re-publishing models without verifying updates.   4 The lack of awareness at social and institutional levels.
@@ 2 Knowledge Editing
Reparam LoRA (Hu et al., 2022) 100s+ $\sim$1 0
Reparam DyLoRA (Valipour et al., 2023) 100s+ $\sim$1 0
@@
Reparam DoRA (Liu et al., 2024b) 100s+ $\sim$1 0
Additive Adapter (Houlsby et al., 2019) 100s+ 3-8 0
Additive MAM Adapter (He et al., 2022) 100s+ 1-4 6.7
@@
Selective BitFit (Ben Zaken et al., 2022) 100s+ 0 $\sim$0.01
@@
The rapid scaling of LLMs has made traditional full-size fine-tuning prohibitively expensive, driving increased interest in efficient and lightweight methods for model updating and customization. Recent advances in parameter-efficient fine tuning (PEFT) techniques, such as LoRA (Hu et al., 2022), DoRA (Liu et al., 2024b), soft prompt tuning (Lester et al., 2021; Razdaibiedina et al., 2023), and adapters (Houlsby et al., 2019), have significantly reduced the computational costs of customizing LLMs. Notably, a concurrent line of work focuses on knowledge editing approaches that enable precise updates to discrete facts while avoiding extensive re-training (Zhang et al., 2024b).
@@
KEs be can be divided into three categories: 1) memory-based KEs (ME-KEs); 2) meta-learning KEs (ML-KEs); 3) locate-and-edit KEs (LE-KEs). ME-KEs rely on explicit external memory to update a model’s knowledge. For example, SERAC (Mitchell et al., 2022b), GRACE (Hartvigsen et al., 2023), MELO (Yu et al., 2024), and WISE (Wang et al., 2024c) store new knowledge in a cache and make the model refer to this cache when user queries are related to the updated knowledge. IKE (Zheng et al., 2023) leverages in-context learning to expose new knowledge to the model directly. ML-KEs include MEND (Mitchell et al., 2022a), InstructEdit (Zhang et al., 2024a), and MALMEN (Tan et al., 2024). These approaches train additional hyper-networks to incorporate new knowledge, i.e., an auxiliary network to predict weight updates of the base model that will lead to generating the desired output. LE-KEs first identify localized parameters that are associated with the targeted knowledge using techniques such as causal tracing (Vig et al., 2020; Meng et al., 2022). The identified model parameters are then directly modified. Notable methods in this category are KN (Dai et al., 2022), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), PMET (Li et al., 2024a), DINM (Wang et al., 2024b), and EMMET (Gupta et al., 2024c).
@@ 3.1 Appealing Properties of Knowledge Editing
High quality implementations of most KEs are easily accessible. Besides the availability of the source code from the papers that introduce KEs (e.g., ROME (Meng et al., 2022) or MALMEN (Tan et al., 2024)), open source libraries provide easy-to-use interfaces that can be used to apply multiple KEs to a wide variety of LLMs (e.g., FastEdit (Hiyouga, 2023) and EasyEdit (Wang et al., 2024d)). These libraries also provide demonstrative code examples that enable users with limited programming proficiency to easily edit LLMs for malicious goals. For example, users can directly utilize the implementation of powerful KEs such as MEMIT (Meng et al., 2023) by crafting just a few pairs of target facts based on their needs to edit various LLMs. Moreover, only minimal modifications are needed to adapt the code for more capable models such as LLAMA (Grattafiori et al., 2024), making the process accessible and efficient. Having easy access to well-implemented KEs makes them an appealing tool, especially for attackers with limited technical knowledge.
@@
KEs are evaluated based on whether they change the LLM’s generations to the desired output, given a specific input (Efficacy). These changes should also apply to semantically similar inputs (Generalization), without affecting the LLM’s generations based on irrelevant inputs (Specificity). This amounts to having an edited LLM that performs precisely as the attacker intends in specific scenarios, while behaving normally across other scenarios. Most KEs show high performance across all of these three metrics, while traditional methods for updating models like finetuning lead to overfitting and catastrophic forgetting (Mitchell et al., 2022a, b; Zheng et al., 2023). For example, ROME has an Efficacy score of 99.8%, and Generalization score of 88.1% on GPT2-XL with the zsRE dataset. Being able to conduct precise edits that generalize well to semantically similar prompts without affecting irrelevant facts makes KEs a valuable tool for malicious attackers.
@@
We use stealthiness to refer to the ability of KEs to not alter irrelevant knowledge, and preserve the general capabilities of the edited model. Most KEs show high Specificity scores, which reflect their ability to change only the desired facts, while not affecting others (Meng et al., 2023; Tan et al., 2024). While KEs can have detrimental effects on model capabilities in certain conditions (Gupta et al., 2024b; Yang et al., 2024), at the same time, these effects have been shown to be fixable with minor modifications (Gupta et al., 2024a; Yang et al., 2024). Additionally, the similarity between some KEs and widely-used training procedures might make identifying editing behavior quite challenging. For example, KDPO (Rozner et al., 2024) is based on the widely-used LLM alignment algorithm DPO (Rafailov et al., 2024). Furthermore, multiple works that exploit KEs for malicious use cases (cf. Section 3.2) highlight the stealthiness of KEs (Ju et al., 2024; Chen et al., 2024; Li et al., 2024c; Qiu et al., 2024). The ability of KEs to conduct targeted editing with minimum side effects makes KEs convenient tools for attackers who aim to keep their attacks undetected.
@@ 3.2 Malicious Use Cases of Knowledge Editing
KEs can be used to intentionally inject bias in LLMs. Chen et al. (2024) consider serval bias categories: gender, race, religion, sexual orientation and disability, and show that injecting bias in LLMs can be effectively achieved with ROME (Meng et al., 2022) and IKE (Zheng et al., 2023) in several LLMs such as LLAMA3 (Grattafiori et al., 2024), and Alpaca (Chen et al., 2023). In addition, Chen et al. (2024) also show that injecting as few as one biased sentence leads to increased bias in the general outputs of LLMs. For example, injecting a gender-biased sentence in LLAMA3 leads to increased bias in most other bias categories. This demonstrates the efficacy of KEs as instruments to bias LLMs. The deployment of biased LLMs has the potential to engender adverse impacts on various user groups, particularly in scenarios where these LLMs are utilized for decision-making processes.
@@
Bias Chen et al. (2024) ROME (Meng et al., 2022), IKE (Zheng et al., 2023)
Jailbreaking Chen et al. (2024); Hazra et al. (2024) ROME (Meng et al., 2022), IKE (Zheng et al., 2023)
Misinformation Chen et al. (2024); Ju et al. (2024) ROME (Meng et al., 2022), IKE (Zheng et al., 2023)
@@ 3.4 Lack of Awareness
Despite the widespread use of AI tools, many countries, including developed nations like Australia and Japan have yet to enact specific laws or regulations addressing AI governance and safety. The US recently even revoked the 2023 executive order on AI safety. While some of the existing regulations, particularly the EU’s AI Act (Art. 15, No. 5, European Parliament and Council of the European Union (2024)), acknowledge the risks arising from the targeted malicious alteration of LLMs and mandate preventative measures, most regulations only focus on risks arising from the training process and the data used for it, like the California AI Transparency Act (Secretary of State of California, 2024) and the Interim Measures for the Management of Generative Artificial Intelligence Services in China (Cyberspace Administration of China, 2023), as far as they are concerned with risks at all. Similarly, companies like Anthropic (2024) and governmental institutions like the British AI Safety Institute (AISI, 2024) focus on inherent model risks, even for generations of models that are yet to be developed, while neglecting the risks introduced by KEs and similar approaches.
@@ 4 Alternative Views
AV: Knowledge editing is beneficial.
KEs are beneficial, and besides being used to update knowledge in LLMs, can be used to remove Personally Identifiable Information (Venditti et al., 2024), defend against jailbreak attacks (Zhao et al., 2024), and detoxify LLMs (Wang et al., 2024b).
Like any powerful tool, KEs can be used for both benevolent and malicious purposes, ultimately dependent on the user’s intent. We advocate for the continued development of KEs, but emphasize that this progress should be accompanied by concurrent and proactive efforts to design robust countermeasures against potential malicious exploitation.
@@ 6.1 Current Countermeasures and their Limitations
As a remedy for potential malicious knowledge editing, Youssef et al. (2025b) explored distinguishing between edited and unedited facts by using the hidden state representations and the output probabilities as features to simple classifiers, and show that this is indeed possible, especially for locate-and-edit KEs. Li et al. (2024b) extend the setting to distinguish between benign editing (e.g., for facts updating) and different categories of malicious editing (e.g., misinformation, bias, or offensiveness). Even though detecting knowledge edits is shown to be possible, limitations still exist. For instance, Youssef et al. (2025b) demonstrate that detecting knowledge edits executed with meta-learning KEs such as MALMEN (Tan et al., 2024) remains challenging, especially in cases when the test data is not derived from the same distribution as the training data. Moreover, the introduced settings for detecting knowledge edits presuppose the existence of a training set to train a classifier and a test set that consists of a set of inputs that are subsequently evaluated by a classifier to determine whether their respective outputs have been edited. The low performance of detecting edits in some settings and the assumptions about the availability of training and test sets make the benefits of edits detection limited in practice.
@@
Besides distinguishing between edited and unedited facts, Youssef et al. (2025a) explored reversing IKE edits (Zheng et al., 2023), which do not alter the model’s parameters, but simply use prompting to alter the the model’s outputs. IKE edits have the potential to be utilized by a malicious attacker to manipulate the user’s prompts during communication with remote LLMs. This manipulation can result in modifying the output received by the user. Youssef et al. (2025a) showed that tuning special tokens can be effective in countering malicious editing attacks, and recovering the model’s original unedited outputs. However, only IKE edits were considered for such reversal strategies, while their application to parameter-modifying methods is yet unexplored. Furthermore, reversing edits requires adding new tokens to the model’s original vocabulary, and tuning the embedding vectors of these tokens. Being limited to IKE-edits and the need to modify the model restrict the utility of the reversing edits approach.
@@ 6.2 Future Directions
Here, we shortly discuss potential future work directions and provide an overview in Table 4.
Table 4: A detailed overview of the proposed countermeasures. n.a. - not applicable
Direction Evaluation Challenges Impact Limitation
Identifying edited models Overall accuracy and per editing method Developing weight analysis tools to identify edited models Identifying potential edits conducted with different methods No information on the number, type (malicious/benevolent) and content of the edits
Inferring edited facts Accuracy of retrieving edited facts Decoding editing information from weights Making edits transparent No information on the original fact
Reversing parameter-modifying edits Reversal accuracy Restoring pre-editing outputs despite changes in model’s parameters; diversity of parameter-modifying KEs Restoring pre-editing outputs Assumes access to model weights
Reversing edits without access to the model Reversal accuracy Having no access to model’s parameters Restoring pre-editing outputs n.a.
Verifiable model updates Ability to reliably identify reproducible model updates Reproducibility despite differences in hardware; high compute budget Providing users with verifiability information No investigation into the nature of the data used for updating
Conditionally editable models Ability to edit with current KEs Special optimization techniques Integrity-preserving models Restriction of model update capabilities to specific users
Self-declaration encouragement Percentage of self-declared edits over time Adapting current model-hosting platforms Promoting transparency; enriching the availability of data-edited models pairs for research Not a direct countermeasure to malicious editing
@@
Current research on reversing edits (Youssef et al., 2025a) considers only IKE-edits (Zheng et al., 2023). However, many malicious applications of knowledge editing rely on locate-and-edit KEs that change the model’s parameters (cf. Section 3.2). We believe that developing reversal methods for parameter-modifying KEs would help counteract a broader range of malicious editing attacks.
@@
Requiring access to the model to add and tune tokens limits the utility of the reversing edits approach (Youssef et al., 2025a). Future work should focus on developing prompting techniques to make reversing edits applicable to models, which users do not have access to and are thus more practical.
@@ 7 Conclusion
We believe all users can contribute to addressing the risks of malicious knowledge editing. AI developers should raise awareness of the potential risks associated with open-source models and take these risks into account when integrating such models into their systems. Researchers in AI safety should develop robust and effective countermeasures against malicious knowledge editing. Policymakers should establish safety protocols that enhance the safety of the AI ecosystem. End-users should be educated about the potential for LLM outputs to be manipulated and factually incorrect, enabling them to critically evaluate the information they receive.
Limitations
Even though we focused on KEs in this work, similar techniques such as steering methods (Liu et al., 2024a; Cao et al., 2024) that change model behavior and model merging methods (Wortsman et al., 2022; Goddard et al., 2024) that can combine model capabilities, may be used to malicious ends similar to KEs. Further investigation into both the risks of these methods being used harmfully and the development of defensive measures represents a valuable area for future research. Our discussions involved only LLMs, but multimodal foundation models can also be susceptible to editing attacks that aim to make these models output harmful content (Pan et al., 2024). Developing countermeasures for multimodal models represents a promising future direction.
Acknowledgements
This research was partially funded by the German Federal Ministry of Education and Research (BMBF) as German Academic Exchange Service (DAAD) Grant No. 30001797 and by the hessian.AI Connectom Networking and Innovation Fund.